(c) List the equilibrium solutions of the differential equation y = (y2 – 1) arctan(y)
Answer #1

For equilibrium solutions, put y'=0

Then, (y+1)(y-1)arctan(y)=0

Which implies y+1=0,y-1=0, arctan (y) =0

Thus, equilibrium solutions are y=-1,1 and 0